an oxide of iron is 77.73% iron by mass; the rest is oxygen. Determine the empirical formula of this compound.

enter a formula
Answer #1

Empirical formula of a chemical compound may be defined as the formula which gives the simplest positive integer ratio of atoms present in a compound.

From the available data given in the question, Percentage of iron= 77.73
Let assume that the total mass is 100 grams
So, the Percentage of oxygen is = 100 - 77.73
= 22.27
We know that molar mass of iron is =55.85
and molar mass of oxygen is = 16

Now we have to calculate the relative mass of iron and oxygen in oxide of iron.

so,
Relative mass of iron in oxide of iron is = Percentage of iron by mass / Atomic mass of iron
= 77.73 / 55.85
= 1.39
Relative mass of oxygen in oxide of iron is = Percentage of oxygen by mass / Atomic mass of oxygen
= 22.27 / 16
=1.39

So,
The simplest molar ratio of iron and oxygen is = 1.39 : 1.39
= 1: 1

So,the Empirical Formula of the compound is FeO